# REPrune: Channel Pruning via Kernel Representative Selection

### 27 Feb 2024 v1

@@ A.1 Preliminary: Ward’s Linkage Distance
A.2 Training Settings
@@
ImageNet dataset consists of 1.27M training images and 50k validation images over 1K classes. The COCO-2017 dataset is divided into train2017, 118k images, and val2017 containing 5k images with 80 object categories. The COCO-2014 provides 83k training images and 41k validation images.
@@
We train ResNet models on the ImageNet dataset using NVIDIA training strategies (Shen et al. 2022), DeepLearningExamples, leveraging mixed precision. Unless otherwise specified, we use default parameters $T_{prune}=180$ and $\Delta T=2$ epochs. Here, $T_{prune}$ indicates the epoch at which channel pruning and regrowth are scheduled, and $\Delta T$ determines the interval to perform the pruning of channels. For data augmentation, we solely adopt the strategies (He et al. 2016), which include random resized cropping to 224$\times$224, random horizontal flipping, and normalization. All models are trained with the same seed. For training ResNet and MobileNetV2 (Sandler et al. 2018) on CIFAR-10, we follow the strategies used for ImageNet but adjust the weight decay to 5e-4 and the batch size to 256. We train the model for 160 epochs without mixed precision and set $T_{prune}=96$.
@@ A.3 Supplementary Experiment
A.4 Instance Segmentation
We evaluate REPrune to the Mask R-CNN, leveraging the COCO-2014 dataset. Mask R-CNN requires the concurrent optimization of three distinct loss functions: $\ell_{cls}$, $\ell_{box}$, and $\ell_{mask}$. This multi-task training presents a challenging scenario for pruning the model’s backbone due to its generalization. Hence, we targeted a 30% reduction in FLOPs for the backbone and examined the results based on a key metric: the Average Precision (AP) for object detection and instance segmentation. According to Table 12, REPrune demonstrates a minimal impact on performance, with a decrease of 0.8 mAP in object detection and 0.4 mAP in instance segmentation. The results still present the efficacy of REPrune in accelerating a more demanding computer vision task.
A.5 Pruning Details
@@
This model is built with multiple bottleneck blocks. Fig.6 illustrates that the bottleneck block also adds the input feature map passing through the skip-connection and the output feature map of the last $1\times 1$ convolution layer within the block. Along with the basic block, if all the convolutional layers are pruned in a bottleneck block, the problem of not performing the addition occurs because the output feature map of the third $1\times 1$ convolution does not match the number of channels with the input feature map of the bottleneck block. In this paper, we perform REPrune only on the first two convolutional layers.

### 08 Mar 2024 v3

@@ A.1 Preliminary: Ward’s Linkage Distance
A.2 Coverage Rate Formula
To quantify the degree of cluster coverage rates after each training epoch, as illustrated in Figure 3, we introduce a metric from Eq. 6, denoted as $\mathcal{R}^{l}_{cov}$ for the $l$-th layer:
$\mathcal{R}^{l}_{cov}=\sum\nolimits_{\mathcal{F}^{l}_{r}\in\tilde{\mathcal{F}}^{l}}\sum\nolimits_{\kappa^{l}_{r,j}\in\mathcal{F}^{l}_{r}}\sum\nolimits_{j\in\mathcal{J}}S(\kappa^{l}_{r,j})/|U^{l}|.$ (14)
This metric calculates the total coverage score $S(\cdot)$ for selected kernels relative to the maximum possible coverage score $|U^{l}|$ across all layers $L$ in a CNN model.
A.3 Training Settings
@@
ImageNet dataset consists of 1.27M training images and 50k validation images over 1K classes. The COCO-2017 dataset is divided into train2017, 118k images, and val2017, which contains 5k images with 80 object categories. The COCO-2014 provides 83k training images and 41k validation images.
@@
We train ResNet models on the ImageNet dataset using NVIDIA training strategies (Shen et al. 2022), DeepLearningExamples, leveraging mixed precision. Unless otherwise specified, we use default parameters $T_{prune}=180$ and $\Delta T=2$ epochs. Here, $T_{prune}$ indicates the epoch at which channel pruning and regrowth are scheduled, and $\Delta T$ determines the interval to perform the pruning of channels. For data augmentation, we solely adopt the strategies (He et al. 2016), which include random resized cropping to 224$\times$224, random horizontal flipping, and normalization. All models are trained with the same seed. For training ResNet and MobileNetV2 on CIFAR-10, we follow the strategies used for ImageNet but adjust the weight decay to 5e-4 and the batch size to 256. We train the model for 160 epochs without mixed precision and set $T_{prune}=96$.
@@ A.4 Supplementary Experiment
A.5 Instance Segmentation
We evaluate REPrune to the Mask R-CNN, leveraging the COCO-2014 dataset. Mask R-CNN requires the concurrent optimization of three distinct loss functions: $\ell_{cls}$, $\ell_{box}$, and $\ell_{mask}$. This multi-task training shows a challenging scenario for pruning the model’s backbone due to its generalization. Hence, we targeted a 30% reduction in FLOPs for the backbone and examined the results based on a key metric: the Average Precision (AP) for object detection and instance segmentation. According to Table 12, REPrune demonstrates a minimal impact on performance, with a decrease of 0.8 mAP in object detection and 0.4 mAP in instance segmentation. The results still present the efficacy of REPrune in accelerating a more demanding computer vision task.
A.6 Pruning Details
@@
This model is built with multiple bottleneck blocks. Figure 6 illustrates that the bottleneck block also adds the input feature map passing through the skip-connection and the output feature map of the last $1\times 1$ convolution layer within the block. Along with the basic block, if all the convolutional layers are pruned in a bottleneck block, the problem of not performing the addition occurs because the output feature map of the third $1\times 1$ convolution does not match the number of channels with the input feature map of the bottleneck block. In this paper, we perform REPrune only on the first two convolutional layers.